hartnn (hartnn):

\(u=e^{(x+y)}+\log (x^3+y^3-x^2 y-xy^2) \) is this an homogeneous function ? how would i prove it ? if it is, is the order 3 ?

OpenStudy (perl):

what is the definition of homogenous function

hartnn (hartnn):

the total exponent of each term should be same

ganeshie8 (ganeshie8):

for homogeneious function we prove \(\large u(tx,ty) = t^ku(x,y)\)

hartnn (hartnn):

\(u=e^{t(x+y)}+\log (x^3+y^3-x^2 y-xy^2)+3 \log t\)

hartnn (hartnn):

the original qustion is State and Prove Euler^' s Theorem for homogeneous functions in two variables and hence find the value of x^2 (∂^2 u)/(∂x^2 ) +2xy (∂^2 u)/∂x∂y+y^2 (∂^2 u)/(∂y^2 )+x ∂u/∂x+y ∂u/∂y for u=e^(x+y)+log (x^3+y^3-x^2 y-xy^2) i did get x^2 (∂^2 u)/(∂x^2 ) +2xy (∂^2 u)/∂x∂y+y^2 (∂^2 u)/(∂y^2 )+x ∂u/∂x+y ∂u/∂y = n^2 u now how do i use this theorem for 'u'

hartnn (hartnn):

is it even homogeneous ?

ganeshie8 (ganeshie8):

yeah it doesn't look homogeneous but we can't say anything unless we prove/disprove it. but from your partial derivatives expression gives the feeling that u is a homogeneous function right ?

hartnn (hartnn):

right, the theorem is asked for homogeneous functions and we are asked to find that value for u ...so i guess they are implying u must be homogeneous

hartnn (hartnn):

if not, should we do something like u = e^a + log b where a ans b are homogeneous , for sure...

ganeshie8 (ganeshie8):

and use chainrule+euler theorem is it ? that looks like a good idea!

hartnn (hartnn):

but euler's will be applicable for a and b only ...

hartnn (hartnn):

\( \partial u/ \partial x = e^a \partial a/ \partial x + 1/b \partial b/ \partial x\) something like this ? won't it become much complicated ...espicially with the 2nd order partial derivative finding...

hartnn (hartnn):

\(u_{xx} = e^a a_{xx} + e^a (a_x)^2 + 1/b b_{xx} -1/b^2 (b_x)^2\) such steps will show up....

hartnn (hartnn):

not sure what to do of terms like (a_x)^2 and (b_x)^2 i hope those get cancelled while finding u_xy

hartnn (hartnn):

\( u_x = e^a a_x + 1/b b_x \) \(u_{xy} = e^a a_y a_x + e_a a_{xy} + 1/b b_{xy} -1/b^2 b_x b_y\)

ganeshie8 (ganeshie8):

but thats just working the partials, not using euler's theorem

hartnn (hartnn):

will be using eurler's for a and b won't calculate a_x, b_x, a_xx, b_xx and so on...

hartnn (hartnn):

still that doesn't work out terms like a_x ^2 and a_x a_y will create problems...

hartnn (hartnn):

any other approach....

Miracrown (miracrown):

homogeneous just means that it's set equal to zero, usually I'm guessing the differential equation is what they mean

Miracrown (miracrown):

What I know about Euler's Theorem is that it's from Number Theory must be some other Euler theorem than from the one I know well, here's what I know about the Euler Theorem: https://en.wikipedia.org/wiki/Euler's_theorem

Miracrown (miracrown):

but I'm guessing that we need a different one here

hartnn (hartnn):

right, this euler theorem is different it states x^2 (∂^2 u)/(∂x^2 ) +2xy (∂^2 u)/∂x∂y+y^2 (∂^2 u)/(∂y^2 )+x ∂u/∂x+y ∂u/∂y = n^2 u for the homogeneous function u of order 'n' that i have already proved

ganeshie8 (ganeshie8):

Oh my! Euler's theorem from number theory would have been fun :D

Miracrown (miracrown):

Yes, indeed.

ganeshie8 (ganeshie8):

I still wonder if that log is in the bottom or exponent

ganeshie8 (ganeshie8):

nvm either way its same level of difficulty >.<

Miracrown (miracrown):

the statement of the Theorem is that whole page that I linked to above ^ and the proof could take a while... let's see where to start

hartnn (hartnn):

i did the proof already!

Miracrown (miracrown):

ok, so they're asking us to restrict it to two variables, that helps

Miracrown (miracrown):

Can I see it?

Miracrown (miracrown):

I don't even see it stated anywhere

hartnn (hartnn):

sure ,just a sec

Miracrown (miracrown):

Wow, okay. Haha. That looks way longer that what I was about to do.

Miracrown (miracrown):

than*

Miracrown (miracrown):

I'm trying to see how the two compare, one sec...

Miracrown (miracrown):

That just seems like a way different approach, but it seems to work. it just relies on a u substitution I can put the other one there as well, just for reference as another approach, I guess... it works for the general theorem, not just for two variables. This one seems specific to two variables - it's also shorter

hartnn (hartnn):

true, lets think of how to use that theorem to find the required expression,,,

Miracrown (miracrown):

http://prntscr.com/4qpdey

Miracrown (miracrown):

Ok, so that's how I would state the theorem and the proof. A bit different, but it accomplishes the same thing I'm sure. ^

Miracrown (miracrown):

So, let's see... both yours and mine wind up with the same formula actually we just need to find what the degree is, the n oh, maybe we just leave it as n... for now at least, but here's the thing all we really need to do is take the second derivative of that statement in the proof. and we'll have something close to what we need.

hartnn (hartnn):

same formula? i am getting n^2 f with that formula, you will be getting 2^n f right ?

Miracrown (miracrown):

ah, I see what they do, it's a bit different. That comes later ^ you're looking at the solution to the whole thing... I'm just talking about the first bold line - no squared terms there

Miracrown (miracrown):

:o I see... yours is better for what we're doing here mine is time based, but yours is based more on x and y derivatives Hmmmm

Miracrown (miracrown):

yeah, it's the same thing though lol the notation just has me thinking too hard

Miracrown (miracrown):

So, at this step, we actually want to take x-derivatives and y-derivatives of both sides that way the n's and t's are constants... now it makes more sense

Miracrown (miracrown):

Solution to the problem: Taking the partial derivative with respect to x on both sides of the previous equation gives us: x d^2f /dx^2 + df/dx + y df/dydx = n df/dx

Miracrown (miracrown):

those should all be partial derivatives ^

Miracrown (miracrown):

now do the same thing but with a y-derivative

Miracrown (miracrown):

x d^2f /dxdy + df/dy + y d^2f/dydx = n df/dy

hartnn (hartnn):

what are we trying to do there ? Lets use 'u' please....are you trying to prove anything here ?? My aim as of now is th find the value of that big expression for 'u' using the theorem i proved...

Miracrown (miracrown):

yeah, ok... you can get lost in all these lines pretty easily

Miracrown (miracrown):

well, we should scrap all of my work then

hartnn (hartnn):

and concentrate on 'u'

Miracrown (miracrown):

we've already done the proof, but I guess we need to do it a different way to have the u in it

Miracrown (miracrown):

seems like you have the solution already I'll just be following that to make sure I get the notation you want otherwise, without that solution, I'd probably do it a slightly different way, on my own.

Miracrown (miracrown):

well, let me just ask you this then what part of the solution that you have that you don't understand fully already?

ganeshie8 (ganeshie8):

seems this question is repelling people in MSE too, nobody is answering http://math.stackexchange.com/questions/947807/if-u-exy-ln-x3y3-x2y-xy2-find-the-value-of

hartnn (hartnn):

i understood everything i did in that proof, but look at the question, there's the 2nd part...

Miracrown (miracrown):

because all we have to do is to multiply n^2 by all of what u is and that's what they're after... that's what the theorem is telling us we can do the proof and the work is long and drawn out but the punchline is to just multiply u by n^2 to get what that entire expression is equal to

ganeshie8 (ganeshie8):

\(u\) is not homogeneous so thats where we're stuck

hartnn (hartnn):

i did get e^(x+y)+ln(x+y)+2ln(x−y) but working out partials is not an option, i think! the question clearly says, "hence" find the value of ...

hartnn (hartnn):

so i need to use the theorem i proved...but how is the question...

Miracrown (miracrown):

then the other stuff, up to the second bold line that's all setup for the form they gave us...since we're dealing with second derivatives they just too both partial derivatives of the form from the theorem and added them together... that's how they get all that stuff on the second bold line which just happens to have the same form as what they are asking for and, due to all the work on that page, winds up being equal to just n^2 times u no matter what u is

Miracrown (miracrown):

see, it seems like this problem assumes a different definition for homogeneous as well no mention of t at all I even have that in my definition of homogeneous

hartnn (hartnn):

but thats true only if u is homogeneous if it is, and if i can prove, then the answer would just be n^2 u, where n is the order of u

Miracrown (miracrown):

right, so do we have to show that this u is homogeneous, and find the order?

Miracrown (miracrown):

that first line where they introduce u for the first time that must be their definition of homogeneous; which is pretty different from mine I don't doubt that they are similar

Miracrown (miracrown):

I think that's what has me thrown off

Miracrown (miracrown):

I see now, I can just take that as a definition (which I can't find anywhere else, by the way) and then the rest seems to make sense now, let me pick it up where you left off by changing variables on u

Miracrown (miracrown):

ah, yeah, I knew something seemed off a bit where did you get this solution? The notation seems a bit wrong, actually in a way They treat x there the way all the other things I'm looking at treat a scalar like the t I was using let me keep digging around, this y/x business has me confused from what I'm looking at, it should be y/t and x/t with the scalar t on the bottom but trying to make it with one variable over the other doesn't seem to agree with the other things I'm seeing

hartnn (hartnn):

u = f(a,b) a and b are functions of x and y u_x in terms of a_x, b_x ?? (it will have a_y, b_y terms also na ?)

hartnn (hartnn):

u_xy, u_xx, u_yy in terms of corresponding 'a's and 'b's ...doesn't seems easy...

Miracrown (miracrown):

I'm just scanning a doc real quick, to make sure i do this corresponding change of variables right

Miracrown (miracrown):

so, there's a change of variables where v = y/x

Miracrown (miracrown):